OpenStudy (anonymous):

What are the intercepts of the equation? Graph the equation. -3x + 4y = -12 i can graph it but i cant find the y and x intercepts from the equation.

OpenStudy (anonymous):

i know its either A. x-intercept: (–3, 0) y-intercept: (0, 4) B. x-intercept: (4, 0) y-intercept: (0, –3)

OpenStudy (radar):

When you substitute 0 for x, y=-3, so that intercept is (0, -3)

OpenStudy (radar):

When you substitute 0 for y getting 0=(3/4)(x)-3 then (3/4) (x)=3 multiplying both sides by 4/3 gives you x=4 for a value of (4,0) Well one is correct, good work.

OpenStudy (radar):

Made error above.I would first express the equation in the slope intercept form y=(3/4)(x) - 3

OpenStudy (anonymous):

thank you!!